(a) Find radius of curvature of curve:

x2+ xy + y2= 4 at point (–2, 0)


1
Expert's answer
2020-06-29T19:14:32-0400

To find Radius of curvature


formula: R= "\frac{(1+y'(x))^{2/3}}{\left | y''(x)) \right |}"

Solution :


y'(x)= "\frac{-(2x+y)}{x+2y}"

At (-2,0), y'(x)= -(-4+0)/(-2)=-2


Y"(x)= [(x+2y)(-2-y')+(2x+y)(1+2y')]/(x+2y)2= (3xy'-3y)/(x+2y)2

At (-2,0), y''(x)= -6)(-2)/4 = 3


Hence substituting in formula , R= "\frac{(1-2)^{2/3}}{\left | 3 \right |}" =-1/3